Show that x inverse is not equal to 0 and is unique


1
Expert's answer
2021-01-26T03:40:53-0500

Here we have "x" "\in\R" (which is invertible).


Now, let "a\in\R" be the multiplicative inverse of "x" . So, by definition we get



"a\cdot x=1_\R (identity\>element\>in\>\R)\\
\Rightarrow a\cdot x=1\\
\Rightarrow x=a^{-1}"

Now, let us assume "a=x^{-1}=0" , so we get

So, we have


"x^{-1}=0\\
\Rightarrow x=\frac{1}{0}=undefined \> in \> \R"

So, we have a contradiction, which arises from our assumption that "a=0" , so we get "a\neq0" .



Now, let us assume "b\in\R" , such that "b" is the multiplicative inverse of "x".


So "b=x^{-1}"


But, we already know that "a=x^{-1}" , so


"\therefore a =b=x^{-1}"


Hence the multiplicative inverse of "x" is unique.